OpenStudy (anonymous):

I need to solve this, then graph it. y=x^-6x+8

OpenStudy (anonymous):

x^2-6x+8 or x^-6x+8 ?

OpenStudy (anonymous):

its a parabola

OpenStudy (anonymous):

I asked if \[x ^{2}-6x+8\] or \[x ^{-6x}+8\] ?

OpenStudy (anonymous):

x squared

OpenStudy (anonymous):

Im trying to figure out how to solve that parabola in standard form

OpenStudy (radar):

y= (x-4)(x-2) those are your zero points for y

OpenStudy (anonymous):

And to graph it , you may need vertex ,for f(x)=ax^2+bx+c , vertex is (-b/2a,f(-b/2a))

OpenStudy (anonymous):

you should write your equation in a complete square form, that's: \[y=(x-3)^2-1\] the vertex is (3,0)

OpenStudy (anonymous):

ok how to solve : to find x the x interecepts and y intercept

OpenStudy (anonymous):

x intercepts and y intercepts

OpenStudy (anonymous):

To find x intercepts , suppose y=0 , y=(x-4)(x-2) so (x-4)(x-2)=0 , x intercepts are 4 and 2 To find y intercept , plug in 0 to x so -4*-2=8

OpenStudy (anonymous):

to find y-int plug x=0, and to find x-int plug y=0. you will get y-int (0,8).. and two x-int (4,0), (2,0)

OpenStudy (anonymous):

I cannot do that because x and x sqaured arent like terms?